# TextureCrop: Enhancing Synthetic Image Detection through Texture-based Cropping

### v2

@@ Abstract
Generative AI technologies produce increasingly realistic imagery, which, despite its potential for creative applications, can also be misused to produce misleading and harmful content. This renders Synthetic Image Detection (SID) methods essential for identifying AI-generated content online. State-of-the-art SID methods typically resize or center-crop input images due to architectural or computational constraints, which hampers the detection of artifacts that appear in high-resolution images. To address this limitation, we propose TextureCrop, an image pre-processing component that can be plugged in any pre-trained SID model to improve its performance. By focusing on high-frequency image parts where generative artifacts are prevalent, TextureCrop enhances SID performance with manageable memory requirements. Experimental results demonstrate a consistent improvement in AUC across various detectors by 6.1% compared to center cropping and by 15% compared to resizing, across high-resolution images from the Forensynths, Synthbuster and TWIGMA datasets.
@@ 1 Introduction
In recent years, the research field of synthetic media generation has seen tremendous growth, as attested by the emergence of numerous state-of-the-art methods [23, 53, 16, 51, 55, 13]. Generative Adversarial Networks (GANs) [19] and Diffusion models [22] have revolutionized the field of synthetic image generation by enabling the creation and editing of images in very high quality and fidelity. While these technologies offer immense creative potential, they also present serious risks for misuse, such as spreading disinformation, facilitating impersonation and fraudulent activities and carrying out foreign interference campaigns. The quality of generated images has reached a level where even the most careful observers can be deceived [34], posing a significant challenge for the majority of Internet users. Moreover, the diversity of fake image sources complicates the detection process, necessitating the development of robust and universal detection methods.
@@
To address this issue, we introduce TextureCrop, an easily adaptable pre-processing SID component that targets high-frequency image patches encapsulating fine details and textures crucial for distinguishing between synthetic and real images. The proposed method, leverages a sliding window technique to systematically analyze image patches and adopts a sampling mechanism that retains patches exhibiting significant texture information while discarding the rest. We conduct an extensive comparison among standard pre-processing techniques during inference to showcase the effectiveness of the proposed method. Experimental results on high-resolution images demonstrate substantial improvements in detection metrics. Specifically, TextureCrop demonstrates an increase in balanced accuracy ranging from 3.4% to 12.1%, an improvement in area under the ROC curve ranging from 6.1% to 14.9%, and an enhancement in average precision ranging from 3.3% to 18.2% across a variety of state-of-the-art SID detectors on more than 10,000 high-resolution real and generated images from the Forensynths [52], Synthbuster [3], TWIGMA [9], RAISE [15] and OpenImages [32] datasets.
Figure 1: Example of the crops created from a 1792x1024 image by using (a) resizing, (b) center cropping, and (c) ten cropping and (d) texture cropping. Images (a) and (b) have the same dimensions ($224\times 224$) as the 10 crops shown in (c) and (d).
@@ 3.1 Problem Formulation
The goal of SID methods is to determine whether an image, $\mathrm{I}\in\mathbb{R}^{C\times W\times H}$, is generated or real, where $C$, $H$, and $W$ denote the number of channels (in most cases 3), height and width of the image, respectively. In the typical formulation, a single image $\mathrm{I}$ (or a batch of images), is fed into the model $f$, resulting in a SID score $s\in[0,1]$, indicating the likelihood of the image being synthetic:
$s=f(\mathrm{I})$ (1)
3.2 Overview
Figure 2: Overview of the TextureCrop Pipeline.
The majority of SID methods often assume that a resized or cropped version of an image is processed, which may result in significant loss of information. Few detectors [20, 11] process the whole image in its native resolution, but in that case processing time and memory requirements increase proportionally to the image’s size. TextureCrop addresses this issue by first dividing image $\mathrm{I}$ into a set of smaller crops $\{\mathrm{C}_{i}\}_{i=1}^{N_{I}}\in\mathbb{R}^{N_{I}\times C\times W^{c}\times H^{c}}$ using a sliding window approach, where $\mathrm{C}_{i}$ represents the i-th crop with dimensions $C\times W^{c}\times H^{c}$ and $N_{I}$ being the number of crops generated for image $I$, which depends on the dimensions of the original image. Subsequently, the crops are ranked according to their texture information (cf. Section 3.3), and the top $\nu$ crops are selected for further analysis. Each crop $\mathrm{C}_{i}$ is processed through the same detection model $f$, returning a SID score $\mathrm{s}_{i}$. The final SID score $s$ is computed by aggregating the individual scores from the selected crops using an aggregation function $\mathfrak{A}$:
@@ 3.3 Selecting texture-rich image crops
Employing a sliding window approach with user-defined stride $\sigma$ and window size $\omega$, where $\omega$ = 224 to align with standard input dimensions for pretrained deep learning models, TextureCrop systematically traverses the input image $\mathrm{I}$ and extracts the corresponding crops $\{\mathrm{C}_{i}\}_{i=1}^{N_{I}}$ similarly to convolving the image. More complex methods, such as Region Proposal Networks (RPNs) [45], were also tested; however, they showed no improvement in detection accuracy despite their significantly increased processing time.
@@ 3.4 Parameters
This subsection outlines the key parameters $\sigma,\omega,\nu,\mu$ used in TextureCrop. Each of these parameters controls different aspects of the cropping behavior. In this section, we will mention the proposed values for these parameters, while their detailed evaluation and experimentation with alternative values will be explored in Sec. 5.2.
@@ 3.4.4 Texture Selection Measure ($\mu$)
While GHE is the best performing measure used in TextureCrop, other alternatives can also be employed, including measures such as Local Entropy (LE), Standard Deviation (SD), Inverse Autocorrelation ($\text{AC}^{-1}$) and Texture Diversity (TD) can capture different aspects of texture information. We explored these alternative measures in the ablations section (see Section 5.2.4) to evaluate their impact on performance.
@@ 3.5 Aggregation method
Our method proposes the processing of a pre-defined number of crops per image, creating the need for an aggregation method to consolidate the scores from individual crops into a single prediction. The best performing aggregation method is the logits’ $\{s_{i}\}_{i=1}^{\nu}$ average:
@@
4 Experimental Setup
4.1 Image Pre-processing Approaches for SID
4.2 Datasets
We evaluate our method using high-resolution data sourced from three datasets: Forensynths [52], Synthbuster [3], and TWIGMA [9]. After filtering out images with resolutions lower than $1024\times 1024$, we collected a total of 2,380 real and generated images from the Forensynths dataset, 3,082 generated images from the Synthbuster dataset, and 1,000 uncompressed and unprocessed high-resolution real images from the RAISE dataset [15]. Additionally, we included a subset of 5,000 generated images from the TWIGMA dataset in our evaluation, along with 1,093 real images from the OpenImages [32] dataset. Table 1 presents an overview of these datasets.
@@ 4.3 SID Detectors
We include the following state-of-the-art detectors in our study: GramNet [33] that leverages global image texture representations, CNNDetect [52] a ResNet50 with blurring and compression augmentation (with two different probabilities of the image being augmented: 0.1 and 0.5), GANID [20] a ResNet50 with intense augmentation that avoids down-sampling in the first layer (trained on ProGAN or StyleGAN2 data), DMID [11], which proposes the fusion of the logits of two models same as described before but one trained on GAN and the other on diffusion data, UnivFD [38] that uses linear probing on features extracted from CLIP’s ViT-L/14 image encoder [17], RINE [31] that uses the image representations extracted by intermediate layers of CLIP’s image encoder (trained on 4 object classes of ProGAN data and on Latent Diffusion Model (LDM) data and PatchCraft [57], which exploits the inter-pixel correlation contrast between rich and poor texture regions within an image. To conduct our experiments, we used the SIDBench framework [48].
@@ 5 Results
5.1 Detection Gains
On average across all detectors and datasets, TextureCrop exhibits an increase of 12.1% increase in BA, 18.2% in AP and 14.9% in AUC compared to Resizing. Compared to CenterCrop, it offers a 4.3% increase in BA, 6.9% in AP and 6.1% in AUC across the 10 methods tested (see Table 2). Compared to TenCrop, it exhibits a 3.4% increase in BA, 3.3% in AP and 6.5% in AUC. The results demonstrate a high degree of consistency across all datasets except for the DALL-E 3 images from Synthbuster, where Resizing seems to lead to slightly better results. These findings highlight the substantial improvements in SID performance achieved with TextureCrop across various metrics and models.
@@ 5.2.3 Number of Crops
We examine the relationship between the number of crops $\nu$, and the effectiveness of TextureCrop. Results are presented in Figures 3(d), 3(e) and 3(f) and reveal a noticeable decrease in performance when only one crop is selected compared to selecting 5 or 10 crops. While the results for 5 and 10 crops are similar, there is a slight performance increase with 10 crops. Despite this increase, processing 10 crops for a high resolution image is not computationally intensive. As such, the associated computational overhead remains modest, justifying the slight performance improvement observed with 10 crops.
@@ 5.2.5 Crop Selection Strategy
Specifically, it shows an average 6.4% decrease in BA, a 10.7% decrease in AP, and a 14.9% decrease in AUC. This highlights that regions with low histogram entropy may not adequately capture important features for effective detection. These differences in performance are illustrated in Figure 6.

### v3

@@ Abstract
Generative AI technologies produce increasingly realistic imagery, which, despite its potential for creative applications, can also be misused to produce misleading and harmful content. This renders Synthetic Image Detection (SID) methods essential for identifying AI-generated content online. State-of-the-art SID methods typically resize or center-crop input images due to architectural or computational constraints, which hampers the detection of artifacts that appear in high-resolution images. To address this limitation, we propose TextureCrop, an image pre-processing component that can be plugged in any pre-trained SID model to improve its performance. By focusing on high-frequency image parts where generative artifacts are prevalent, TextureCrop enhances SID performance with manageable memory requirements. Experimental results demonstrate a consistent improvement in AUC across various detectors by 6.1% compared to center cropping and by 15% compared to resizing, across high-resolution images from the Forensynths, Synthbuster and TWIGMA datasets. Code available at https://github.com/mever-team/texture-crop.
@@ 1 Introduction
Figure 1: Example of the crops created from a 1792x1024 image by using (a) resizing, (b) center cropping, (c) ten cropping and (d) texture cropping. Images (a) and (b) have the same dimensions ($224\times 224$) as the 10 crops shown in (c) and (d).
In recent years, the research field of synthetic media generation has seen tremendous growth, as attested by the emergence of numerous state-of-the-art methods [23, 54, 16, 52, 56, 13]. Generative Adversarial Networks (GANs) [19] and Diffusion models [22] have revolutionized the field of synthetic image generation by enabling the creation and editing of images in very high quality and fidelity. While these technologies offer immense creative potential, they also present serious risks for misuse, such as spreading disinformation, facilitating impersonation and fraudulent activities, compromising biometric security systems and carrying out foreign interference campaigns. The quality of generated images has reached a level where even the most careful observers can be deceived [34, 39], posing a significant challenge for the majority of Internet users. Moreover, the diversity of fake image sources complicates the detection process, necessitating the development of robust and universal detection methods.
@@
To address this issue, we introduce TextureCrop, an easily adaptable pre-processing SID component specifically designed for high-resolution images. The proposed method targets high-frequency image patches that encapsulate fine details and textures crucial for distinguishing between synthetic and real images. It leverages a sliding window technique to systematically analyze image patches and adopts a sampling mechanism that retains patches exhibiting significant texture information while discarding the rest. We conduct an extensive comparison among standard pre-processing techniques during inference to showcase the effectiveness of the proposed method. Experimental results on high-resolution images demonstrate substantial improvements in detection metrics. Specifically, TextureCrop demonstrates an increase in balanced accuracy ranging from 3.4% to 12.1%, an improvement in area under the ROC curve ranging from 6.1% to 14.9%, and an enhancement in average precision ranging from 3.3% to 18.2% across a variety of state-of-the-art SID detectors on more than 10,000 high-resolution real and generated images from the Forensynths [53], Synthbuster [3], TWIGMA [9], RAISE [15] and OpenImages [32] datasets.
$s=f(\mathrm{I})$ (1)
1.1 Overview
Figure 2: Overview of the TextureCrop Pipeline.
@@ 3.1 Problem Formulation
The goal of SID methods is to determine whether an image, $\mathrm{I}\in\mathbb{R}^{K\times W\times H}$, is generated or real, where $C$, $H$, and $W$ denote the number of channels (in most cases 3), height and width of the image, respectively. In the typical formulation, a single image $\mathrm{I}$ (or a batch of images), is fed into the model $f$, resulting in a SID score $s\in[0,1]$, indicating the likelihood of the image being synthetic:
The majority of SID methods often assume that a resized or cropped version of an image is processed, which may result in significant loss of information. Few detectors [20, 11] process the whole image in its native resolution, but in that case processing time and memory requirements increase proportionally to the image’s size. TextureCrop addresses this issue by first dividing image $\mathrm{I}$ into a set of smaller crops $\{\mathrm{C}_{i}\}_{i=1}^{N_{I}}\in\mathbb{R}^{N_{I}\times K\times W^{c}\times H^{c}}$ using a sliding window approach, where $\mathrm{C}_{i}$ represents the i-th crop with dimensions $C\times W^{c}\times H^{c}$ and $N_{I}$ being the number of crops generated for image $I$, which depends on the dimensions of the original image. Subsequently, the crops are ranked according to their texture information (cf. Section 3.2), and the top $\nu$ crops are selected for further analysis. Each crop $\mathrm{C}_{i}$ is processed through the same detection model $f$, returning a SID score $\mathrm{s}_{i}$. The final SID score $s$ is computed by aggregating the individual scores from the selected crops using an aggregation function $\mathfrak{A}$:
@@ 3.2 Selecting texture-rich image crops
Employing a sliding window approach with user-defined stride $\sigma$ and window size $\omega$, where $\omega$ = 224 to align with standard input dimensions for pretrained deep learning models, TextureCrop systematically traverses the input image $\mathrm{I}$ and extracts the corresponding crops $\{\mathrm{C}_{i}\}_{i=1}^{N_{I}}$ similarly to convolving the image. For images with dimensions that are not exact multiples of the window size, the method ensures complete coverage by including additional overlapping crops. Specifically, when the sliding window reaches the edge of the image along either the width or height, and the remaining area is less than the window size, the algorithm selects crops starting from the last valid position while overlapping with previously selected regions. More complex methods, such as Region Proposal Networks (RPNs) [46], were also tested; however, they showed no improvement in detection accuracy despite their significantly increased processing time.
@@ 3.3 Parameters
This subsection outlines the key parameters $\sigma,\omega,\nu,\mu$ used in TextureCrop. Each of these parameters controls different aspects of the cropping behavior. In this section, we mention the proposed values for these parameters, while their detailed evaluation and experimentation with alternative values is presented in Sec. 5.2.
@@ 3.3.4 Texture Selection Measure ($\mu$)
While GHE is the best performing measure used in TextureCrop, other alternatives can also be employed, including measures such as Local Entropy (LE), Standard Deviation (SD), Inverse Autocorrelation ($\text{AC}^{-1}$) and Texture Diversity (TD), which can capture different aspects of texture information. We explored these alternative measures in the ablations section (see Section 5.2.4) to evaluate their impact on performance.
@@ 3.4 Aggregation method
Our method proposes the processing of a pre-defined number of crops per image, creating the need for an aggregation method to consolidate the scores from individual crops into a single prediction. The best performing aggregation method was found to be the logits’ $\{s_{i}\}_{i=1}^{\nu}$ average:
@@
4 Experimental Setup
4.1 Image Pre-processing Approaches for SID
4.2 Datasets
We evaluate our method using high-resolution images sourced from three datasets: Forensynths [53], Synthbuster [3], and TWIGMA [9]. After filtering out images with resolutions lower than $1024\times 1024$, we collected a total of 2,380 real and generated images from the Forensynths dataset, 3,082 generated images from the Synthbuster dataset, and all 1,000 uncompressed and unprocessed high-resolution real images from the RAISE dataset [15]. Additionally, we included a subset of 5,000 generated images from the TWIGMA dataset in our evaluation, along with 1,093 high-resolution real images from the OpenImages [32] dataset, selected also based on the criterion of having a resolution of $1024\times 1024$ pixels or greater. Table 1 presents an overview of these datasets.
@@ 4.3 SID Detectors
We include the following state-of-the-art pre-trained detectors in our study: GramNet [33] that leverages global image texture representations, CNNDetect [53] a ResNet50 with blurring and compression augmentation (with two different probabilities of the image being augmented: 0.1 and 0.5), GANID [20] a ResNet50 with intense augmentation that avoids down-sampling in the first layer (trained on ProGAN or StyleGAN2 data), DMID [11], which proposes the fusion of the logits of two models same as described before but one trained on GAN and the other on diffusion data, UnivFD [38] that uses linear probing on features extracted from CLIP’s ViT-L/14 image encoder [17], RINE [31] that uses the image representations extracted by intermediate layers of CLIP’s image encoder (trained on 4 object classes of ProGAN data and on Latent Diffusion Model (LDM) data and PatchCraft [58], which exploits the inter-pixel correlation contrast between rich and poor texture regions within an image. To conduct our experiments, we used the SIDBench framework [49]. The checkpoints for all detectors were obtained from the original repositories of the respective papers.
@@ 5 Results
5.1 Performance Gains
@@
On average across all detectors and datasets, TextureCrop exhibits an increase of 12.1% increase in BA, 18.2% in AP and 14.9% in AUC compared to Resizing. Compared to CenterCrop, it offers a 4.3% increase in BA, 6.9% in AP and 6.1% in AUC across the 10 methods tested (see Table 2). Compared to TenCrop, it exhibits a 3.4% increase in BA, 3.3% in AP and 6.5% in AUC. Among the methods tested, PatchCraft demonstrates the largest gains with a 10.00% increase in BA, 9.14% in AP, and 9.64% in AUC, highlighting its strong performance with TextureCrop. On the other hand, RINE LDM shows the smallest improvements, with a slight decrease of 1.32% in BA, a small increase of 0.21% in AP, and a negligible decrease of 0.15% in AUC.
CenterCrop and Resizing consistently display worse performance on average across the datasets compared to TextureCrop, across all methods. In contrast, TenCrop is the only method that occasionally outperforms TextureCrop, but even in these cases, it is only by a slight margin. The results demonstrate a high degree of consistency across all datasets except for the DALL-E 3 images from Synthbuster, where Resizing seems to lead to slightly better results.
It is also worth noting that, compared to the methods that process the entire image, TextureCrop significantly reduces computational overhead. For example, for a $2048\times 2048$ image, TextureCrop reduces processing time from 1.05s to 0.63s and lowers memory usage from 17 GB to 2.08 GB. Additionally, it addresses the limitations of methods that process the entire image, which are often constrained by memory and computational restrictions, as these methods can only handle images up to a certain size.
@@ 5.2.3 Number of Crops
We examine the relationship between the number of crops $\nu$, and the effectiveness of TextureCrop. Results are presented in Figures 3(d), 3(e) and 3(f) and reveal a noticeable decrease in performance when only one crop is selected compared to selecting 5 or 10 crops. While the results for 5 and 10 crops are similar, there is a slight performance increase with 10 crops. Despite this increase, processing 10 crops for a high resolution image is not computationally intensive. As such, the associated computational overhead remains modest, and is justified by the performance improvement observed with 10 crops.
@@ 5.2.5 Crop Selection Strategy
Specifically, it shows an average 6.4% decrease in BA, a 10.7% decrease in AP, and a 14.9% decrease in AUC. This highlights that regions with low histogram entropy do not adequately capture important features for effective detection. These differences in performance are illustrated in Figure 6.
@@ 6 Conclusion
7 Acknowledgments
This work is partially funded by the Horizon 2020 European project vera.ai under grant agreement no. 101070093.